The overall population for Region A is 90 million people. The labor force contains 44 million people, 35 million people are employed, and 9 million are unemployed. What is the unemployment rate? Round your answer to the nearest whole number.Suppose that a baby boom occurs between the years 2025-2030, similar to the baby boom following World War II. What do you expect to happen to the labor force participation rate in the year 2090-2100, holding all else constant? The labor force participation rate will O increase as the baby boomers will continue to work while in the influx of young workers will drive up the participation rate decrease since the baby boomers will opt for more leisure. O not change since every retiree will have a new worker replace them. O increase since the baby boomers will opt for more leisure.
Answer #1

Unemployment rate = unemployed/labor force * 100 = 9/44*100 = 0.2045 * 100 = 20.45%

Answer 2: Decrease as the baby boomers will opt for more lesisure.

Explanation: People born in 2025-2030 will be aged enough to retire or choose to work less in 2090-2100.